OpenStudy (anonymous):

Which statement best expresses the revenue (r) for selling (n) tickets at $5.00 each?

OpenStudy (anonymous):

r = 5.00n r + n = 5.00 r – n = 5.00 r=n/5.00

OpenStudy (baru):

what do you think it is?

OpenStudy (anonymous):

i dont know thats why i aksed you all